If the side of a square is decreased by 3 meters, the area is decreased by 45 square meters. Find the length of the original square.
Answer by vksarvepalli(154)

let the side of original square be x meters
so area will be x^2 sq.mt
now given If the side of a square is decreased by 3 meters, the area is decreased by 45 square meters
so when the side is x-3 the area is (x-3)^2 that is given equal to x^2-45
so (x-3)^2 = x^2-45
=> x^2-6x+9 = x^2-45
=>6x=54
so x=9
so the side of the original square is 9 meters in length